OpenStudy (anonymous):

The height of a cylinder is 3 times larger than the radius. Find the volume of the cylinder as a function of the height

OpenStudy (anonymous):

\[\pi h^3/9\]

OpenStudy (anonymous):

the multiple choice answers are V = 3πh3 V = 1/9 πh3 V = 9πh3 V = 1/3 πh3